OpenStudy (anonymous):

Find the additive inverse (opposite) of: - 4.7

OpenStudy (anonymous):

The additive inverse is the number which when added will equal the identity (normally 0) so the additive inverse of -4.7 is positive 4.7.

OpenStudy (anonymous):

Can you please help me with another one and Thank you so much. 3 ( ( - 5 ) + 7 (( 19 - 7 ) / 4 )

OpenStudy (anonymous):

I think you are missing a parenthisis, maybe after the 4?... 4))? Double check to see if a parenthesis is extra or if one is missing? Perhaps there is only one ( after the 3?

OpenStudy (anonymous):

3 ( (-5)+7)+((19-7)/4

OpenStudy (anonymous):

There is still either an extra or a missing parenthesis. This makes it difficult to solve without guessing. Check it carefully again. :) My guess is that there should only be one ( before the number 19 - in which case the answer is 3(2) + (12)/4 = 6 + 3 = 9.

OpenStudy (anonymous):

3 ( (-5 )+ 7 ) + ((19-7)/4) I'm so sorry this is the right way.

OpenStudy (anonymous):

No problem, so the answer will still be 3(2) + (12/4) = 6 + 3 = 9. I kind of assumed. :) You always have to solve the parts in parenthesis first...starting at the innermost and moving out.

OpenStudy (anonymous):

Thank you so much (: